OpenStudy (anonymous):

Which graph models the equation -3x + 2y = 6? What number does it fall on the x axis and y axis

OpenStudy (jim766):

first solve for y, do you know how to do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

do i add 3 to both sides

OpenStudy (jim766):

-3x + 2y = 6 yes +3x +3x 2y = 3x + 6 now what?

OpenStudy (jim766):

what is keeping the y from being alone?

OpenStudy (anonymous):

uhmmm. it has 2 in front of it

OpenStudy (jim766):

right....it is 2times y , do the opposite to get rid of it

OpenStudy (jim766):

divide both sides by 2

OpenStudy (anonymous):

3x + 3

OpenStudy (anonymous):

but on a graph on my x axis will my line cross three and cross the y axis to enter my 2nd quadrant

OpenStudy (jim766):

-3x + 2y = 6 yes +3x +3x 2y = 3x + 6 now what? -- -- -- 2 2 2 y = 3/2x + 3

OpenStudy (jim766):

when you divide both sides by 2, you have to divide everything by the 2

OpenStudy (jim766):

y = mx + b y = 3/2x + 3

OpenStudy (anonymous):

oh ok so when i divide 3 by 2 whats the next time after that

OpenStudy (jim766):

6/ 2 = 3 y = mx + b y = 3/2x + 3 what is the y intercept?

OpenStudy (anonymous):

2x i think

OpenStudy (anonymous):

NO 3

OpenStudy (anonymous):

so itll be 2 on the x axis and cross and fall on the 3 on the y axis if im correct

OpenStudy (jim766):

yes 3 is the y intercept, that is where the line crosses the y axis set y = to 0 and solve for x to see where it crosses the x axis -3x + 2y = 6 -3x + 2(0) = 6 solve for x

OpenStudy (anonymous):

ok when i multiply and my 2 is gone do i add 3x to both sides or.. 6 im assuming 3 but then i get um 3=6

OpenStudy (anonymous):

and now i guess if it is 3=6 i um divide 3 by both sides which leads to the conclusion of 3/6 = 2 right ?

OpenStudy (jim766):

divide both sides by -3 6/-3 = -2

OpenStudy (jim766):

the line crosses the x axis at -2

OpenStudy (anonymous):

thanks jim can you help me with with other problem

OpenStudy (anonymous):

A geometric sequence is defined by the recursive formula t1 = 64, tn = tn - 1 2 , where n ∈N and n > 1. The sequence is

OpenStudy (jim766):

t1 = 64, tn = tn -2 , is that right?

OpenStudy (anonymous):

i have to find the next 6 terms starting with 64

OpenStudy (anonymous):

A) 64, 16, 8, 4, 2, 1, ... Eliminate B) 64, 32, 16, 8, 4, 2, ... C) 64, 128, 256, 512, 1024, 2048, ... D) 128, 256, 512, 1024, 2048, 4096, ... these are the choices